OpenStudy (anonymous):

Differentiate the following function. h (θ) = 7 sec(θ) + 4 eθ cot(θ)

OpenStudy (anonymous):

this is what i got...\[7(\sec \theta \tan \theta)+4(e ^{\theta} \cot \theta -\csc ^{2}\theta)\]

OpenStudy (anonymous):

\[dh/d \theta=7\sec \theta \tan \theta-4e^\theta\csc \theta \cot \theta-4e^\theta \csc \theta\]

OpenStudy (anonymous):

oops my last part was wrong

OpenStudy (anonymous):

whats wrong with mine?

OpenStudy (anonymous):

your looks fine...one sec...lemme double check

OpenStudy (anonymous):

webassign is telling me its wrong

OpenStudy (anonymous):

it is...i see what u did wrong

OpenStudy (anonymous):

the last part is -e^(theta)csc^2(theta))

OpenStudy (anonymous):

by chain rule

OpenStudy (anonymous):

so its 4(-e^theta csc^2(theta))??

OpenStudy (anonymous):

7(secθtanθ)+4(eθcotθ−(*(e^(theta))*)csc2θ

OpenStudy (anonymous):

take all the parenthesis out...that starred area is what u needed to add...

OpenStudy (anonymous):

u get it?